OpenStudy (anonymous):

Please Solve DE: (x+1)dy/dx=10-y

OpenStudy (amistre64):

split the variable?

OpenStudy (anonymous):

Im not very good with diff Q's

OpenStudy (anonymous):

this is a normal differential equation. you still need help with it?

OpenStudy (anonymous):

yes please. If you know how

OpenStudy (amistre64):

split it into 2 parts; dy and dx; then integrate both sides

OpenStudy (anonymous):

Integrating isnt hard, just starting the problem is.

OpenStudy (amistre64):

\[ \frac{(x+1)dy}{dx}=10-y\] \[ \frac{dy}{10-y}=\frac{dx}{x+1}\]

OpenStudy (anonymous):

yep .. i was writing it out but its simple as that. separable differential equation

OpenStudy (amistre64):

give the left a -1/-1 and run with it :)

OpenStudy (anonymous):

wouldnt the dx be below x+1 or no

OpenStudy (anonymous):

Im not sure if thats the final answer

OpenStudy (anonymous):

you need x terms with dx and y terms with dy in a separable equation.

OpenStudy (anonymous):

x+1)dy = (10-y)dx , integrating both sides: (x+1)y = (10-y)x + c , [c=constant], xy+y=10x-xy + c , 2xy +y =10x +c